If B=7i -3j+2k and D = 4i +5j 3k. The angle between B and D is
1
Expert's answer
2020-12-04T11:14:29-0500

"|B|=\sqrt{7^2+3^2+2^2}=7.874"


"|D|=\sqrt{4^2+5^2+3^2}=7.071"


"B\cdot D =7\cdot4+3\cdot5+2\cdot3=49"


"\cos\alpha=\frac{B\cdot D}{|B||D|}=\frac{49}{7.874\cdot 7.071}=0.88\to \alpha=28.4"